6. (4 marks) The moment generating function (mgf) of a random variable X is given by m(t)-e2 (a) Use the mgf to find the mean and variance of X (b) What is the probability that X-2?
Answer #1

a)

here first derivative of mgf m'(t)=(d/dt)*m(t) =(d/dt)*e2t =2e2t

therefore mean E(X)=m'(0)=2e2*0 =2

second derivative of mgf m''(t)=(d/dt)*m'(t)=(d/dt)*2e2t =4e2t

hence E(X2) =m''(0)=4e2*0 =4

hence Variance =E(X2)-(E(X))2 =4-22 =0

b)

as mgf =px1*etx1+px2*etx2+px3*etx3+...

comaring it with given mgf:

P(X=2)=1